# Learning Equivariant Maps with Variational Quantum Circuits

### v1

@@ IV.4 Example: $S_{6}$ Symmetry
V Discussions
@@
We have presented a novel approach to learning equivariant maps between two representations of finite groups through the use of variational quantum circuits. By parameterizing the embedding of the data space into Hilbert space, and training it to respect the inherent symmetries of the system, we have demonstrated a method for capturing and preserving group symmetries in quantum computations. Our exploration of various finite groups highlights the versatility and effectiveness of this approach, and each example showcases the capacity of variational quantum circuits to learn and respect different symmetries, thereby ensuring that equivariant maps are accurately learned.
@@
It should be noted that the choice of architecture in the embedding can have a profound influence on the ability of a model to learn equivariant maps. Indeed, an architecture which is too limited in scope may be incapable of representing an equivariant map between two particular representations. This can happen if there are too few parameters in the model or the gate set used is not expressive enough. On the other hand, an architecture that is too broad in scope might hinder the ability of the model to train. This can happen, for example, if there are too many parameters in the model or the gate set is overly expressive.
Overall, the results presented here contribute to the theme of geometric quantum machine learning by showing that an important aspect of this theory, the embedding, can itself be learned. Moreover, our results give a method for testing whether two representations of a finite group are isomorphic, and when this is not the case, it allows us to produce a measure of just how similar they are.
Acknowledgements.
@@
The authors acknowledge support from the DoD Smart Scholarship and OUSD(R&E) SMART SEED grant. This document has been approved for public release, Distribution A. Distribution is unlimited.

### v2

@@ III Learning Equivariant Maps
IV Learning Intertwining Maps
We call a linear equivariant map between two representations an intertwining map. Such maps preserve both the linear vector space structure and the symmetry structure afforded by the representation. When the intertwining map is also a bijection, we call it an isomorphism of representations, and the representations are called equivalent. In representation theory, we seek to classify all irreducible representations according to this equivalence relation.
A limitation of the approach in Section III is that an embedding of the data space into Hilbert space is necessarily nonlinear, and so such an equivariant map cannot transfer the linear structure of $\mathcal{R}$ onto $\mathcal{H}$. Indeed, if the embedding were linear, then $\mathcal{E}(2x)=2\mathcal{E}(x)$. But if $\mathcal{E}(x)\in\mathcal{U}(\mathcal{H})$, then $2\mathcal{E}(x)\not\in\mathcal{U}(\mathcal{H})$, a contradiction. Fortunately, there is a way around this issue.
Suppose that we are given a dataset with some symmetry described by the representation $V:G\to GL(\mathcal{R})$. Let us denote by $\mathcal{E}:\mathcal{R}\to\mathcal{U}(\mathcal{H})$ the amplitude embedding, so that $T:\mathcal{R}\to\mathcal{H}$ defined by $T(x)=\mathcal{E}(x)\ket{\vec{0}}$ is the map sending each data element to its corresponding quantum state. While the amplitude embedding $\mathcal{E}$ is not linear, its unnormalized counterpart is. Indeed, the map ^$\hat{\mathcal{E}}:\mathcal{R}\to\mathcal{U}(\mathcal{H})$ defined by ^$\hat{\mathcal{E}}(x)=\|x\|\mathcal{E}(x)$ produces the linear operator which performs an unnormalized amplitude embedding, and so the Hilbert space level linear operator that we are looking for is $\hat{T}:\mathcal{R}\to\mathcal{H}$ defined by $\hat{T}=\hat{\mathcal{E}}(x)\ket{\vec{0}}$. The trouble is that ^$\hat{\mathcal{E}}(x)$ is not a unitary operator, and so we cannot implement it in a straightforward way in our quantum circuit. Instead, we resort to the following proposition linking the linear equivariance of $\hat{T}$ to the equivariance of $T$.
Lemma 3.
Let $T(x)=\mathcal{E}(x)\ket{\vec{0}}$ where $\mathcal{E}$ is the normalized amplitude embedding, and let $\hat{T}(x)=\hat{\mathcal{E}}(x)\ket{\vec{0}}$ where $\hat{\mathcal{E}}$ is the unnormalized amplitude embedding. Then $\hat{T}$ is an intertwining map between unitary representations $V$ and $W$ if and only if $T$ is equivariant between the same representations.
Proof.
It is clear that $\hat{T}$ is linear as it maps the vector $x_{0}e_{0}+\cdots+x_{n}e_{n}\in\mathcal{R}$ to $x_{0}\ket{0}+\cdots+x_{n}\ket{n}\in\mathcal{H}$. Let $V$ be the representation on $\mathcal{R}$ and let $W$ be the representation on $\mathcal{H}$. Then for all $g\in G$, we have
$W(g)\hat{T}(x)=\hat{T}(V(g)x)$ (31)
if and only if
$\frac{1}{\|x\|}W(g)\hat{T}(x)=\frac{1}{\|x\|}\hat{T}(V(g)x),$ (32)
which is equivalent to
$W(g)T(x)=\frac{\|V(g)x\|}{\|x\|}T(V(g)x).$ (33)
Thus, if $V$ is unitary, then the assumption that $\hat{T}$ is an intertwining map is equivalent to the condition
$W(g)T(x)=T(V(g)x),$ (34)
which is precisely the condition that $T$ is an equivariant map. ∎
Lemma 3 assumes that the embedding is the usual unparameterized amplitude embedding, and so its linear extension cannot learn an intertwining map between representations. Instead, we insert a parameterized unitary operator which we will use to learn equivariance between the representation induced by $\hat{T}$ and another representation $W$ acting on the same Hilbert space. Let us denote by $A:G\to\mathcal{U}(\mathcal{H})$ the representation induced by $\hat{T}$, which we recall satisfies
$\hat{T}\circ V(g)=A(g)\circ\hat{T}.$ (35)
We begin by performing the amplitude embedding on our data, after which we apply a parameterized unitary operator $U_{\theta}$. We then perform an operator $U_{inv}$ which is $G$-invariant with respect to the $W$ representation and measure with respect to an observable $\mathcal{O}$ that is likewise $G$-invariant. Let us again denote by $h_{\theta}$ the expected value of this circuit. We now claim that when $U_{\theta}$ is an equivariant map between $A$ and $W$, the estimate $h_{\theta}$ is $G$-invariant with respect to $V$.
Proposition 2.
Let $V:G\to\mathcal{U}(\mathcal{R})$ and $W:G\to\mathcal{U}(\mathcal{H})$ be representations. The estimate
$h_{\theta}(x)=\operatorname{Tr}\left[U_{inv}U_{\theta}T(x)(U_{inv}U_{\theta}T(x))^{\dagger}\mathcal{O}\right]$ (36)
is $G$-invariant (i.e. $h_{\theta}(V(g)x)=h_{\theta}(x)$) whenever the unitary $U_{inv}$ and the observable $\mathcal{O}$ are $G$-invariant with respect to $W$ and $U_{\theta}$ is an intertwining map between $W$ and the representation $A$ induced by $\hat{T}$.
Proof.
Indeed, we have
$h_{\theta}(V(g)x)=\operatorname{Tr}\left[U_{inv}U_{\theta}T(V(g)x)\left(U_{inv}U_{\theta}T(V(g)x)\right)^{\dagger}\mathcal{O}\right].$ (37)
Since $A$ is induced by $\hat{T}$, we have by Lemma 3 that $T$ is an equivariant map between $V$ and $A$, and it follows that
$=\operatorname{Tr}\left[U_{inv}U_{\theta}A(g)T(x)\left(U_{inv}U_{\theta}A(g)T(x)\right)^{\dagger}\mathcal{O}\right].$ (38)
Now applying the equivariance of $U_{\theta}$ produces
$=\operatorname{Tr}\left[U_{inv}W(g)U_{\theta}T(x)\left(U_{inv}W(g)U_{\theta}T(x)\right)^{\dagger}\mathcal{O}\right].$ (39)
The $G$-invariance of $U_{inv}$ then gives
$=\operatorname{Tr}\left[W(g)U_{inv}U_{\theta}T(x)\left(W(g)U_{inv}U_{\theta}T(x)\right)^{\dagger}\mathcal{O}\right]$ (40)
$=\operatorname{Tr}\left[W(g)U_{inv}U_{\theta}T(x)(T(x))^{\dagger}U_{\theta}^{\dagger}U_{inv}^{\dagger}(W(g))^{\dagger}\mathcal{O}\right]$ (41)
$=\operatorname{Tr}\left[U_{inv}U_{\theta}T(x)(T(x))^{\dagger}U_{\theta}^{\dagger}U_{inv}^{\dagger}(W(g))^{\dagger}\mathcal{O}W(g)\right],$ (42)
where in the last equality, we have applied the cyclicity of the trace. Now applying the $G$-invariance of the observable $\mathcal{O}$ produces
$=\operatorname{Tr}\left[U_{inv}U_{\theta}T(x)(T(x))^{\dagger}U_{\theta}^{\dagger}U_{inv}^{\dagger}(W(g))^{\dagger}W(g)\mathcal{O}\right]$ (43)
$=\operatorname{Tr}\left[U_{inv}U_{\theta}T(x)(T(x))^{\dagger}U_{\theta}^{\dagger}U_{inv}^{\dagger}\mathcal{O}\right]$ (44)
$=\operatorname{Tr}\left[U_{inv}U_{\theta}T(x)(U_{inv}U_{\theta}T(x))^{\dagger}\mathcal{O}\right]$ (45)
$=h_{\theta}(x).$ (46)
This completes the proof. ∎
To train the equivariance of $U_{\theta}$, we sample from both the data space and the group and minimize the loss function
$L_{g}(x)=(h_{\theta}(V(g)x)-h_{\theta}(x))^{2}$ (47)
just as in Section III, and again optimize using one of the many available optimization routines. Thus, we have a procedure for learning an intertwining map between the representation $A$ induced by $\hat{T}$ and the given representation $W$.
V Conclusion
@@
We have presented a novel approach to learning equivariant maps between representations of finite groups using variational quantum circuits. By parameterizing the embedding of the data space into Hilbert space, and training it to respect the inherent symmetries of the system, we have demonstrated a method for capturing and preserving group symmetries in quantum computations. Our exploration of various finite groups highlights the versatility and effectiveness of this approach, and each example showcases the capacity of variational quantum circuits to learn and respect different symmetries.
@@
It should be noted that the choice of architecture in the embedding can have a profound influence on the ability of a model to learn equivariant maps. Indeed, an architecture which is too limited in scope may be incapable of representing an equivariant map between two particular representations. This can happen if there are too few parameters in the model or the gate set used is not expressive enough. The trivial example is a standard embedding procedure, which is not parameterized at all and therefore cannot learn an equivariant map. On the other hand, an architecture that is too broad in scope might hinder the ability of the model to train. This can happen, for example, if there are too many parameters in the model or the gate set is overly expressive. A well-designed architecture strikes a balance between expressiveness and trainability, avoiding limitations due to insufficient parameters while mitigating the challenges of over-parameterization.
Our findings contribute to the field of geometric quantum machine learning by showing that an important aspect of this theory, the embedding, can itself be learned. Moreover, this work provides a method for measuring the similarity between two representations, paving the way for further advancements in symmetry-based quantum information processing. These results demonstrate the potential for variational quantum circuits to be powerful tools for symmetry-respecting tasks.
Acknowledgements
@@
MLL, ZPB, and ENE acknowledge support from the DoD Smart Scholarship and OUSD(R&E) SMART SEED grant.
This document has been approved for public release, Distribution A. Distribution is unlimited.